# LLM Attention Transplant for Transfer Learning of Tabular Data Across Disparate Domains

### v1

@@ Abstract
Transfer learning of tabular data is non-trivial due to heterogeneity in the feature space across disparate domains. The limited success of traditional deep learning in tabular knowledge transfer can be advanced by leveraging large language models (LLMs). However, the efficacy of LLMs often stagnates for mixed data types structured in tables due to the limitations of text prompts and in-context learning. We propose a lightweight transfer learning framework that fine-tunes an LLM using source tabular data and transplants the LLM’s selective keykey and valuevalue projection weights into a gated feature tokenized transformer (gFTT) built for tabular data. The gFTT model with cross-domain attention is fine-tuned using target tabular data for transfer learning, eliminating the need for shared features, LLM prompt engineering, and large-scale pretrained models. Our experiments using ten pairs of source-target data sets and 12 baselines demonstrate the superiority of the proposed LLM-attention transplant for transfer learning (LATTLE) method over traditional ML models, state-of-the-art deep tabular architectures, and transfer learning models trained on thousands to billions of tabular samples. The proposed attention transfer demonstrates an effective solution to learning relationships between data tables using an LLM in a low-resource learning environment. The source code of the proposed method is shared in  [^1] https://anonymous.4open.science/r/LATTLE—LLM-Tabular-Transfer-Learning-667E
@@ 1.1 Literature review and contributions
Recent studies have optimized LLM weights by fine-tuning [Yang2024_iTabLLM, Dinh2022_LIFT]. Gardner et al. [Gardner2024_TABULA-8B] propose TABULA-8B, a LLaMA-3-based language model fine-tuned on billions of tabular data rows serialized as text. LLM fine-tuning in similar studies uses an Auto-Regressive loss that predicts the next token in the sequence [Gardner2024_TABULA-8B, Dinh2022_LIFT, Zhang2023_LLaMAGTL, Hegselmann2023_TabLLM], which differs from the learning objective of a classifier model. Furthermore, the pretrained TABULA-8B model supports few-shot in-context learning of downstream target data, where the performance gain plateaus after 32 shots in the prompt due to the inherent limitation of ICL. We argue that downstream classification can also involve model fine-tuning to overcome the limitations of ICL. Moreover, their pretrained model learns from eight billion tabular data rows, demanding extensive computing resources, whereas a lightweight transfer learning can be devised with better performance. In this paper, we propose a new LLM Attention Transplant for Transfer LEarning (LATTLE) to address the challenges of cross-domain transfer learning of tabular data using an LLM. We show that transplanting selective attention-related weights from a fine-tuned LLM into a transformer model specifically built for tabular data can yield superior transfer-learning performance without requiring common features, prompt engineering, and large data samples. The following are the main contributions of this paper.
@@
  • Transfer learning learns the relationship between tabular data sets of different domains without any common features.
  • The proposed method performs downstream learning using a native transformer model for tabular data without the need for in-context learning and LLM prompt engineering.
@@ 2.1 Preliminaries
Here, $W_{q}$, $W_{k}$, $W_{v}$} $\in\mathbb{R}^{d_{k}\times d}$ are trainable weight matrices, and $q_{j}$, $k_{j}$, $v_{j}$} $\in\mathbb{R}^{d_{k}}$ are the resulting queryquery, keykey, and valuevalue vectors for feature $j$. The attention of feature $r$ to feature $j$ ($\alpha_{rj}$) is computed using a scaled dot product of the queryquery vector of the $r$ feature and the keykey vector of the $j$ feature.
$\alpha_{rj}=\text{softmax}\left(\frac{q_{r}^{\top}k_{j}}{\sqrt{d_{k}}}\right),\quad Z_{r}=\sum_{j=1}^{m}\alpha_{rj}v_{j}$ (2)
Here, $Z_{r}\in\mathbb{R}^{d_{k}}$ represents the attention-weighted context vector of feature $r$ across all $m$ features. It should be noted that attention weights are directional because $\alpha_{rj}\neq\alpha_{jr}$.
@@
2.2 Fine-tuning LLM with source data
The first step of the proposed transfer learning method is to acquire the information of the source tabular data set in an LLM. Serialization of pairs of feature values of tabular data in a textual description (e.g. “Age is 25. Sex is male.”) is tokenized and encoded, as shown in Figure 1. Because the seminal GPT 3.5 model is closed-source, we use an open source and lightweight version of GPT2 [radford2019_GPT2], known as DistilGPT2 [huggingface_distilgpt2]. DistilGPT2 reduces computational cost by 50% while retaining approximately 97% of the original GPT-2 performance. DistilGPT2 has six attention layers and accepts input sequences of up to 1024 tokens. The tokenizer (Byte Pair Encoding (BPE)) of DistilGPT2 is used to obtain token IDs and the corresponding embedding vectors from pretrained DistilGPT2. These embeddings are used to fine-tune the uppermost layer of a DistilGPT2 model for the classification task related to the source data set.
The embeddings of keykey, queryquery, and valuevalue (Equation 1) are reshaped through a projection layer to be compatible with a gated feature tokenizer transformer (gFTT) with five attention layers. The DistilGPT2-gFTT coupling facilitates text-to-tabular knowledge transfer, which is followed by a fully connected classifier head. Supervised fine-tuning of the DistilGPT2 model using the source tabular data is performed using the cross-entropy loss defined in Equation 3.
@@ 2.2 Fine-tuning LLM with source data
Models: LLM (source), gFTT (target)
@@
Output: $gFTT_{CA}$ finetuned with LLM attention
LLM Fine-tuning: Source Data Set
@@
  $LM_{s} \leftarrow LL$M with gFTT head $(X_{s},y_{s})$ (Eq. 3)
@@
 Weights of $L$ Attention Layers of LLMLLM :
$\{W^{1},...,W^{L}\}$ $\leftarrow$ $LLM_{s}$
$\{W^{L}_{q},W^{L}_{k},W^{L}_{v}\}$ $\leftarrow$ $W^{L}$, the top attention layer,
gFTT Training: Target Data Set
 Initialize gFTT with Attention Layers $l_{0}$, $l_{1}$, …, $l_{4}$ }
Cross-attention: LLM Weight Transplanted in gFTT
  gFTT $W_{k}$, $W_{v}$$_{\{\ell_{0}\}}$ $\leftarrow[W_{k}^{L},W_{v}^{L}]$ Frozen
@@
  $gFTT_{CA}\leftarrow gFTT(X_{t},y_{t})$
@@ 2.3 Proposed LATTLE algorithm
The LLM-based transfer learning method proposed for tabular data is summarized in Algorithm 1 and illustrated in Figure 1. LLM weights of the uppermost attention layer are fine-tuned by the source data set and then transplanted into a state-of-the-art transformer model for tabular data, gated feature tokenized transformer (gFTT) [Wang2022_Transtab]. Specifically, the keykey and valuevalue projection weights ($W^{s}_{k}$ and $W^{s}_{v}$) of the uppermost layer of a fine-tuned LLM are transplanted as frozen weights in the lowest attention layer of the gFTT model. To facilitate this transplant, the gFTT model has attention weights of the same size (768×768) as those in DistilGPT2.
@@ 4 Results
DB 0.822 (0.04) 0.803 (0.03) 0.798 (0.03) 0.818 (0.03) 0.788 (0.04) 0.643 (0.03)
52
AutoML
data
sets
0.834 (0.02)
Open
Tabs
2000
data
sets
0.803 (0.03)
Eight
billion
tabular
rows
0.680*
Distil-
GPT2
0.793 (0.04) CD 0.799 (0.04) 0.653 (0.06) 0.831 (0.04)
@@
DB 0.747 (0.04) 0.720 (0.04) 0.753 (0.03) 0.772 (0.03) 0.754 (0.04) 0.625 (0.05)
52
AutoML
data
sets
0.764 (0.02)
OpenTabs
2000
data
sets
0.734 (0.02)
Eight
billion
tabular
rows
N/R
Distil-
GPT
0.724 (0.03) CD 0.734 (0.04) 0.610 (0.07) 0.753 (0.03)

### v2

# LATTLE: LLM Attention Transplant for Transfer Learning of Tabular Data Across Disparate Domains

@@ Abstract
Transfer learning on tabular data is challenging due to disparate feature spaces across domains, in contrast to the homogeneous structures of image and text. Large language models (LLMs) offer a knowledge base to improve the limited effectiveness of cross-domain transfer learning for tabular data. However, LLM performance often stagnates due to subjective text prompts and the computational limitations of in-context learning. We present a novel language-to-tabular context-learning method that uses attention-specific transformer weights, enabling seamless transfer learning across disparate tabular data sets. The LLM attention transplant mechanism facilitates a domain-agnostic transfer learning, eliminating the need for shared features between tables, LLM prompt engineering, and large-scale pretrained models. Our experiments using ten pairs of disjoint source-target data sets and 12 baseline methods demonstrate the superiority of the proposed LLM-attention transplant for transfer learning (LATTLE) method over traditional ML models, state-of-the-art deep tabular architectures, and models trained on thousands to billions of tabular samples. The proposed cross-domain attention transfer demonstrates an effective solution for adapting LLMs to learning non-text tabular data in a low-resource environment. The source code of the proposed method is shared in  [^1] https://anonymous.4open.science/r/LATTLE—LLM-Tabular-Transfer-Learning-667E
@@ 1.1 Literature review and contributions
Recent studies have optimized LLM weights by fine-tuning [Yang2024_iTabLLM, Dinh2022_LIFT]. Gardner et al. [Gardner2024_TABULA-8B] propose TABULA-8B, a LLaMA-3-based language model fine-tuned on billions of tabular data rows serialized as text. LLM fine-tuning in similar studies uses an Auto-Regressive loss that predicts the next token in the sequence [Gardner2024_TABULA-8B, Dinh2022_LIFT, Zhang2023_LLaMAGTL, Hegselmann2023_TabLLM], which differs from the learning objective of a classifier model. Furthermore, the pretrained TABULA-8B model supports few-shot in-context learning on downstream target data, but performance gains plateau after 32 shots in the prompt due to the inherent limitations of ICL. We argue that downstream classification can also involve model fine-tuning to overcome the limitations of ICL. Moreover, their pretrained model is trained on eight billion tabular data rows, requiring extensive computing resources, whereas a lightweight transfer learning approach can achieve better performance.
1.2 Contributions
In this paper, we propose a new LLM Attention Transplant for Transfer LEarning (LATTLE) to address the challenges of cross-domain tabular data transfer learning using cross-attention with an LLM. The contributions of our paper are as follows.
@@ 1.2 Contributions
  • The LLM is trained using a learning objective designed for tabular data, rather than the standard auto-regressive loss that predicts the next token in a sequence.
  • The context between heterogeneous domains is captured by transplanting attention weights from an LLM, which obviates the requirement for shared features between different tabular datasets.
  • The proposed method performs downstream cross-attention learning using finetuned LLM weights without the need for prompt engineering and in-context learning.
@@ 2.1 Preliminaries
Here, $W_{q}$, $W_{k}$, $W_{v}$} $\in\mathbb{R}^{d_{k}\times d}$ are trainable weight matrices, and $q_{j}$, $k_{j}$, $v_{j}$} $\in\mathbb{R}^{d_{k}}$ are the resulting queryquery, keykey, and valuevalue vectors for feature $j$. The attention of feature $p$ to feature $j$ ($\alpha_{pj}$) is computed using a scaled dot product of the queryquery vector of the $p$ feature and the keykey vector of the $j$ feature.
$\alpha_{pj}=\text{softmax}\left(\frac{q_{p}^{\top}k_{j}}{\sqrt{d_{k}}}\right),\quad Z_{p}=\sum_{j=1}^{m}\alpha_{pj}v_{j}$ (2)
Here, $Z_{p}\in\mathbb{R}^{d_{k}}$ represents the attention-weighted context vector of feature $p$ across all $m$ features. It should be noted that attention weights are directional because $\alpha_{pj}\neq\alpha_{jp}$.
@@
2.2 Fine-tuning LLM using source tabular data
The first step of the proposed transfer learning method is to adapt an LLM to a tabular data learning task. In particular, LLMs are trained using an autoregressive loss to predict the next token in a sentence. Adapting language models to mixed-type tabular data requires a tailored data preprocessing framework. The first step involves serialization of feature value pairs of tabular data in a textual description (e.g. “Age is 25. Sex is male.”), which are then tokenized and encoded, as shown in Figure 1. It is important to note that the seminal GPT‑3.5 model and its successor LLMs are proprietary and not open-source. Although more advanced open-source LLMs can be finetuned with expensive computational resources, they may have already been trained on publicly available data, making them prone to data memorization. Accordingly, we adopt DistilGPT2 [huggingface_distilgpt2], an open-source, lightweight variant of GPT2 [radford2019_GPT2], to demonstrate the effectiveness of language-to-tabular cross-attention learning. DistilGPT2 reduces the computational cost by 50% while retaining approximately 97% of the original performance of GPT2. DistilGPT2 has six attention layers and accepts input sequences of up to 1024 tokens. The tokenizer (Byte Pair Encoding (BPE)) of DistilGPT2 is used to obtain token IDs and the corresponding embedding vectors from pretrained DistilGPT2. To enable LLM-to-tabular context learning, we adopt a state-of-the-art transformer designed for tabular data, the gated feature tokenizer transformer (gFTT)  [Wang2022_Transtab] together with DistilGPT2, forming an end-to-end learning framework. As transformer-based architectures, both gFTT and DistillGPT2 possess trainable projection matrices associated with the keykey, queryquery, and valuevalue representations (see Equation 1). These projection weights are used to streamline language-to-tabular knowledge transfer within the DistilGPT2-gFTT joint framework, which is followed by a classifier head. The top layer of DistilGPT2 is fine-tuned jointly with gFTT using the source tabular data to perform a supervised classification task defined by the cross-entropy loss in Equation 3.
@@ 2.2 Fine-tuning LLM using source tabular data
Models: LLM (Source), gFTT (Head), gFTT (Target)
@@
Output: $gFTT_{CA}$ cross-table transfer via LLM attention
LLM Fine-tuning on Source: Language-to-Tabular Alignment
 LLM-gFTT $\leftarrow$ {LLM (Source), gFTT (Head)}
@@
  $\mathcal{L}^{Source}_{\mathrm{CE}}\leftarrow\mathcal{L}(W^{\mathrm{LLM}}_{r,L},\{W^{gFTT}_{r,\ell}\};X_{s},y_{s})$ (Eq. 6)
  Update LLM and gFTT Attention Weights $\leftarrow$ $W^{LLM}_{r,L}$, $W^{gFTT}_{r,\ell}$ (Eq. 45)
@@
 Weights of $L$ Attention Layers of LLMLLM : $\{W^{1},...,W^{L}\}$
Key-Value Projection Weights from LLM: $\{W^{LLM}_{K,L},W^{LLM}_{V,L}\}$ $\leftarrow$ $W^{LLM}_{r,L}$
gFTT (Target) as $t$ : Source-to-Target Transfer Learning
  $W^{t}_{K,0}\leftarrow W^{LLM}_{K,L};\quad W^{t}_{V,0}\leftarrow W^{LLM}_{V,L}$ (Frozen)
@@
  Cross-attention, $\alpha_{s,t}\leftarrow\text{Attention}(W_{q}^{t},W_{k}^{LLM},f_{t})$ (Eq. 7)
  Attention weighted embedding, $Z_{t}\leftarrow\alpha_{s,t}W_{v}^{LLM}f_{t}$ (Eq. 8)
  $\mathcal{L}^{Target}_{\mathrm{CE}}\leftarrow\ell(W^{t}_{Q,0},\{W^{t}_{r,\ell}\}_{\ell=1}^{4};X_{t},y_{t})$ (Eq. 9)
  Update $\{W^{t}_{Q,0},W^{t}_{r,\ell>0}\}$ minimizing $\mathcal{L}^{Target}_{\mathrm{CE}}$ (Eq. 10)
@@
2.3 Hierarchical attention weight selection
Hierarchical representation learning has been well studied in convolutional neural networks for vision tasks. However, the distribution of knowledge in the attention layers of transformers for tabular data remains underexplored. We use empirical evidence on knowledge distribution from the literature to select and transfer attention weights for computing cross-domain (LLM-to-tabular) attention learning. The LLM trained using source data has finetuned projection weights $W_{Q}^{S}$, $W_{K}^{S}$, and $W_{V}^{S}$ corresponding to the query, key, and value representations, respectively. Among these, the key and value projection weights ($W_{K}^{S}$, $W_{V}^{S}$) encode contextual representations learned from the source tabular data (Equation 1). The literature suggests that the representations learned by a transformer are distributed hierarchically across layers, where higher layers tend to capture richer contextual knowledge [dai2022]. Therefore, we use the key and value projection weights ($W_{K}^{S}$, $W_{V}^{S}$) of the topmost attention layer of LLM for downstream attention transplantation. Similarly, the downstream transformer model learns target tabular data, where the most contextual target knowledge is at the topmost layer. In contrast, the base, or lowest, layer encodes more general knowledge and is typically kept frozen to preserve this information during downstream transfer learning. This approach is empirically supported by [hwang2025_freezinglower], which shows that freezing the bottom layers enhances performance in downstream training. Consequently, the base layer of the downstream transformer model remains frozen with the LLM weights ($W_{K}^{S}$, $W_{V}^{S}$), allowing the upper layers to incorporate generalized LLM context while also learning task-specific knowledge for the target tabular data. The LLM-to-tabular context learning mechanisms are explained in the next section.
@@ 2.4 Proposed LATTLE algorithm
The LLM-based transfer learning method proposed for tabular data is summarized in Algorithm 1 and illustrated in Figure 1. We learn the language-to-tabular context by finetuning a joint LLM-gFTT framework using source tabular data, as mentioned in Section 2.2. The transformer projection weights ($W_{q}$$W_{k}$$W_{v}$) of LLM (topmost layer, $L$) and gFTT (all five layers, $l$) are updated to optimize cross-entropy loss as follows.
$W^{LLM}_{r,L}\leftarrow W^{LLM}_{r,L}-\eta\frac{\partial\mathcal{L}_{\mathrm{CE}}(X_{s})}{\partial W^{LLM}_{r,L}};\quad\forall_{r}\in\{Q,K,V\}$ (4)
$W^{gFTT}_{r,\ell}\;\leftarrow\;W^{gFTT}_{r,\ell}\;-\;\eta\frac{\partial\mathcal{L}_{\mathrm{CE}}(X_{s})}{\partial W^{gFTT}_{r,\ell}};\quad\forall_{\ell}\in\{0,\dots,4\},\ \forall_{r}\in\{Q,K,V\}$ (5)
The projection weights of the transformers in LLM and gFTT are jointly updated to align the LLM with tabular data representations of gFTT optimizing a common cross-entropy loss in Equation 6 using source data ($X_{s}$).
$\mathcal{L}^{Source}_{\mathrm{CE}}=\mathcal{L}_{\mathrm{CE}}\Big[y_{s},f(W^{\mathrm{LLM}}_{r,L},\;\{W^{gFTT}_{r,\ell}\}_{\ell=0}^{4},X_{s})\Big],\forall_{r}\in\{Q,K,V\}$ (6)
The training prepares the projection weights of the LLM for learning the general context of tabular data. The finetuned LLM weights corresponding to keykey and valuevalue representations ($W_{k}$$W_{v}$) contain a general knowledge base suitable for learning tabular data. Therefore, this weight pair is kept frozen at the base of a new downstream gFTT to provide the foundational context of the LLM while updating the upper layers using the target tabular data. In the base layer of the gFTT transformer, the frozen projection weights ($W_{k}^{LLM}$$W_{v}^{LLM}$) of the LLM interact with the unfrozen weight of the queryquery representation ($W_{q}^{t}$). Therefore, cross-domain attention $\alpha_{s,t}$ between the LLM tuned to the source and the target is achieved at the weight level of the transformer, as shown in Equation 7, instead of using keykey, valuevalue, queryquery representations in Equation 2.
$\alpha_{s,t}=\text{softmax}\left(\frac{(W_{q}^{t}f_{t})^{\top}W_{k}^{LLM}f_{t}}{\sqrt{d_{k}}}\right).\quad$ (7)
Here, $f_{t}$ is the embedding of the target tabular data, explained in the next section. In other words, the target queryquery to LLM keykey achieves language-to-tabular attention, which is used to achieve cross-domain tabular transfer learning. Attention weighted ($\alpha_{s,t}$) target data embedding ($Z_{t}$) is obtained following Equation 8.
$Z_{t}=\alpha_{s,t}W_{v}^{LLM}f_{t}$ (8)
The downstream gFTT is optimized for the classification task specific to the target tabular data set.
$\mathcal{L}^{Target}_{\mathrm{CE}}=\mathcal{L}_{\mathrm{CE}}\Big[y_{t},f(W^{t}_{Q,0},\;\{W^{t}_{r,\ell}\}_{\ell=1}^{4},X_{t})\Big],\forall_{r}\in\{Q,K,V\},$ (9)
where, LLM weight transplantation, $W^{t}_{K,0}$ = $W^{LLM}_{K,L}$ and $W^{t}_{V,0}$ = $W^{LLM}_{V,L}$, is performed in downstream with the constraint shown in Equation 10.
$\frac{\partial\mathcal{L}_{\mathrm{CE}}^{Target}}{\partial W^{t}_{K,0}}=0,\frac{\partial\mathcal{L}_{\mathrm{CE}}^{Target}}{\partial W^{t}_{V,0}}=0$ (10)
3 Language-encoded tabular embedding
@@ 5 Results
DB 0.822 (0.04) 0.803 (0.03) 0.798 (0.03) 0.818 (0.03) 0.788 (0.04) 0.643 (0.03) 52 AutoML data sets 0.834 (0.02) Open Tabs 2000 data sets 0.803 (0.03) Eight billion tabular rows 0.680* Distil- GPT2 0.793 (0.04) CD 0.799 (0.04) 0.653 (0.06) 0.831 (0.04)
@@
DB 0.747 (0.04) 0.720 (0.04) 0.753 (0.03) 0.772 (0.03) 0.754 (0.04) 0.625 (0.05) 52 AutoML data sets 0.764 (0.02) OpenTabs 2000 data sets 0.734 (0.02) Eight billion tabular rows N/R Distil- GPT 0.724 (0.03) CD 0.734 (0.04) 0.610 (0.07) 0.753 (0.03)